OpenStudy (gabbyalicorn):

What is the slope of the line? y= -5x-9 slope= ?

OpenStudy (anonymous):

Well your formula is y=mxb right?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

Wait a sec.

OpenStudy (gabbyalicorn):

mk

OpenStudy (anonymous):

Alright your slope is -5 The key to this is... It's the number before the x

OpenStudy (abdullah1995):

y = mx + b M = slope ----------------- y = mx+b y = -5x-9 when comparing the general line equation and your specific equation, we can see that "y" are the same in both, "x" are the same in both, however m=-5 and b = -9. Since 'M" is the slope, your slope in this equation will be -5.

OpenStudy (gabbyalicorn):

so should I do 5x- 9 I'm still confused?

OpenStudy (anonymous):

Okay Here, This might help you y=mx+b Take this equation for an example. y=y m= -5 (SLOPE) b= y INTERCEPT So if you were to graph this it would be... I hope i drew this well enough!

OpenStudy (anonymous):

Yes it took twenty minutes XD

OpenStudy (gabbyalicorn):

so the slopes -5x

OpenStudy (anonymous):

No. The Slope Is Just The NUMBER before x Like... y=-2x+6 Your Slope (The letter M in the Equation Y=MX+B) Would Be...?

OpenStudy (gabbyalicorn):

-5

OpenStudy (gabbyalicorn):

:D?

OpenStudy (gabbyalicorn):

Thanks for helping me

OpenStudy (anonymous):

No... This Equasion I've given you is y=-2x+6 in EVERY one like this it will always be the Number Before the letter x in that equation... Right? So If I were to have one that looked like the one above my slope (letter M) would be the number and the sign after the =